OpenStudy (anonymous):

i really need help with this. derivative of y=cos(2x), then y=?

OpenStudy (nubeer):

d/dx cos (2x) d/dx(2x) -sin(2x) *2

OpenStudy (anonymous):

Use the chain rule: \[y'=\frac{d}{d(2x)}(\cos(2x)) \times \frac{d}{dx}(2x)=-\sin(2x) \times 2=-2\sin(2x)\]